# Diffusive Gibbs Sampling

### v2

@@ 2.2 Choosing the Gaussian Convolution Kernels
The performance of DiGS can be influenced by the hyperparameters $\alpha,\sigma$ in the Gaussian convolution kernel. Intuitively, for a given fixed $\alpha$, a large $\sigma$ will enhance the exploration capability of the sampler. However, this also makes the denoising posterior $p(x|\tilde{x})$ closer to $p(x)$ as illustrated in Equation 11, thereby elevating the complexity of denoising sampling. Similarly, with a fixed $\sigma$, reducing $\alpha$ will bring the modes closer but will also make $\nabla_{x}\log p(x|\tilde{x})$ close to $\nabla_{x}\log p(x)$ as shown in Equation 12, making the denoising sampling challenging. To illustrate the effects of hyperparameters, we apply DiGS to the MoG problem described in Figure 1 with varying values of $\alpha,\sigma$ and show the results in Figures 4(a) and 4(b), respectively. This demonstrates that within a specific range of $\alpha$ and $\sigma$ values, DiGS consistently achieves optimal performance indicated by almost zero MMD. However, the quality of samples degrades when hyperparameters deviate beyond certain thresholds.
Although these results suggest that DiGS is robust to variations in hyperparameters within a certain range, selecting the appropriate parameter range for a specific target problem remains crucial. The optimal range often depends on the support or structure of the target density, which is typically unknown in practice. In the next section, we introduce a multi-level Gaussian convolution schedule to mitigate the need for precise hyperparameter selection.
@@ 2.3 Multi-Level Noise Scheduling
Drawing inspiration from the noise scheduling technique used in diffusion models (Ho et al., 2020; Song & Ermon, 2019; Song et al., 2021; Gao et al., 2018), we propose to use a sequence of Gaussian convolution kernels with $10{<}\alpha_{T}{<}\cdots{<}\alpha_{1}{<}$ 1 and the corresponding variance $\sigma^{2}_{t}=1-\alpha_{t}^{2}$ in each Gaussian convolution kernel $p_{t}(\tilde{x}_{t}|x)=\mathcal{N}(\tilde{x}|\alpha_{t}x,\sigma_{t}^{2}I)$. This approach is commonly known as the variance-preserving (VP) schedule, where $p_{t}(\tilde{x})=\int p_{t}(\tilde{x}_{t}|x)p(x)\mathop{}\!\mathrm{d}{x}$ and $p_{0}(\tilde{x})\equiv p(x)$. As $\alpha_{T}\rightarrow 0$, it follows that $p_{T}(\tilde{x}_{T}|x)\rightarrow\mathcal{N}(0,I)$ and $p_{T}(\tilde{x}_{T})\rightarrow\mathcal{N}(0,I)$, which is independent of the characteristics of the initial distribution. We follow the common sampling procedure in the score-based diffusion model literature (Song & Ermon, 2019, 2020) and propose the following sampling procedure: for each $t$ from $T$ to $1$, we run DiGS to generate a sample, which is used as an initial point in the subsequent time step $t{-}1$. For a given number of noise levels $T$, we apply a simple linear scheduling scheme to determine the intermediate $\alpha_{t},\sigma_{t}$ values. Specifically, given a $\alpha_{1}$ and $\alpha_{T}$, we can determine $\alpha_{t}$ by $\alpha_{t}=\alpha_{T}+(T-t)\Delta_{\alpha}$, where $\Delta_{\alpha}=(\alpha_{1}-\alpha_{T})/(T-1)$ and $\sigma_{t}=\sqrt{1-\alpha_{t}}$.
@@ 3 Comparison to Related Methods
Figure 5: Illustration of a multi-modal target distribution $p(x)$, tempered distribution $p_{\beta}(x)$, and convolved distribution $p(\tilde{x})$.
@@ 3.1 Tempering-Based Sampling
where $\beta\equiv 1/\tau<1$ is the inverse temperature. As $\tau\rightarrow\infty$, the tempered target $p_{\beta\rightarrow 0}(x)$ converges to a flat distribution, which encourages transitions among different modes. Tempering is a key building block to develop state-of-the-art multi-modal sampling methods such as parallel tempering (PT) (Swendsen & Wang, 1986; Geyer & Thompson, 1995) and annealed importance sampling (AIS) (Neal, 2001).
@@
To empirically compare the mode-bridging property of PT and DiGS, we conduct experiments on an extreme problem, “Mixture of Deltas”, which is an MoG with extremely small variance in each Gaussian. Figure 6 shows the comparison of the samples from these two methods. We find that PT is unable to escape the central mode in this extreme setting even with a large temperature $\tau$. On the other hand, DiGS manages to recover all 9 modes, demonstrating its superiority over PT in terms of mode exploration and coverage.
@@
Figure 6: Comparison between parallel tempering (PT) and DiGS on a “Mixture of Deltas” problem in 2D. PT consists of 5 temperatures, while DiGS uses one noise level. Each sampler is initialized at the origin and generates 1,000 samples. Detailed experimential setup can be found in Appendix C.4.
@@ 3.2 Score-Based Diffusion Model
The convolution technique plays a key role in score-based diffusion models (Song & Ermon, 2019; Song et al., 2021; Song & Ermon, 2020; Ho et al., 2020; Sohl-Dickstein et al., 2015; Gao et al., 2020). These models employ a series of convolutions (indexed by $t$) to form a distribution sequence from $p(\tilde{x}_{0})\equiv p(x)$ to $p(\tilde{x}_{T})$. At each time step $t$, the score function at time $\nabla_{\tilde{x}_{t}}\log p(\tilde{x}_{t})$ is learned from samples ${x_{1},\cdots,x_{N}}\sim p(x)$ using denoising score matching (Vincent, 2011). To sample from $p(x)$, ULA and Euler discretization are typically applied in reverse ($t=T\to 0$), using samples from each time step $t$ as the initial point at time $t{-}1$. This scheme has shown state-of-the-art generation quality for complex data like images. However, unlike score-based/diffusion generative models where $\nabla_{\tilde{x}}\log p(\tilde{x})$ is learned directly from samples, our setting only assumes access to an energy function $E(x)$, making the noisy marginal score $\nabla_{\tilde{x}}\log p(\tilde{x})$ intractable, as shown in Equation 9.
@@
Zhang et al. (2023a) proposes a pseudo-Gibbs sampler to sample from the joint $p(x,\tilde{x})=p(x)p(\tilde{x}|x)$ where $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$ and $p(x|\tilde{x})$ is approximated by a full-covariance Gaussian. This approach also requires training data to estimate the score function $\nabla_{\tilde{x}}\log p(\tilde{x})$, which is different from our problem setting.
3.3 Proximal MCMC
@@
The proposed DiGS also belong to the proximal MCMC family, as referenced in Chen et al. (2022); Lee et al. (2021), where Gibbs sampling is conducted between the Gaussian convolution distribution $p(\tilde{x}|x)$ and the denoising distribution $p(x|\tilde{x})$. We highlight several significant improvements over the proximal MCMC method discussed in Chen et al. (2022). First, we employ a score-based sampler, such as ULA or HMC, instead of the rejection sampling in the denoising step, which helps the method scale to higher-dimensional distributions. Second, we utilize an MH scheme to initialize the denoising sampling, demonstrating its importance in achieving correct density allocation in multi-modal situations, as demonstrated in Section 2.1. Third, we emphasize the importance of selecting the Gaussian convolution kernel in Section 2.3 and extend DiGS to a multi-noise setting, eliminating the need to choose Gaussian convolution hyperparameters. These improvements make the method applicable to practical problems.
@@
Entropy-MCMC (Li & Zhang, 2023) is designed to sample from the flat regions within the posterior of the Bayesian neural network $p(\theta|\mathcal{D})$, where $\mathcal{D}$ represents the dataset. Specifically, the approach begins by defining a surrogate distribution using Gaussian convolution:
@@
where $p(\theta_{a}|\theta)\propto\exp\left(-\frac{1}{2\eta}\|\theta_{a}-\theta\|^{2}_{2}\right)$ is a Gaussian kernel aimed at smoothing out the sharp regions in the posterior. Subsequently, SGLD is applied to sample within the joint space of $(\theta,\theta_{a})$. It is crucial to highlight that, while this method is aimed at optimization to seek flat minima in the target distribution $p(\theta|\mathcal{D})$, similar to the approaches described in Chaudhari et al. (2019); Staines & Barber (2012), it is not intended and cannot provide exact samples from the target distribution $p(\theta|\mathcal{D})$. In contrast, DiGS is designed as a valid MCMC sampler that is specifically designed to identify all modes (including both sharp and flat modes) and their corresponding density allocations in the target distribution, thus pursuing a distinct objective.
Figure 7: Computational cost comparison between DiGS and RDMC. The x-axis is the number of energy evaluations (×$\times 10^{6}$) and the y-axis is the MMD between true and generated samples. We use DiGS with a single noise level $(\alpha=1,\sigma=1)$ and run Gibbs sampling for 1-10 sweeps, represented by the 10 blue scatters. The RDMC is experimented with $T\in\{1,2,3,4\}$, represented by the four orange points in the plot. Detailed experimental setup can be found in Appendix C.5.
@@
3.4 Reverse Diffusion Monte Carlo
$\nabla_{x_{t}}\log p(x_{t})$ $=\frac{\alpha\mu(x_{t})-x_{t}}{\sigma_{t}^{2}},$ (19)
where the denoising posterior mean $\mu(x_{t})=\int xp(x|x_{t})\approx\frac{1}{K}\sum_{k=1}^{K}x_{0}^{(t,k)}$ is approximated by the Monte Carlo method, with samples $x_{0}^{(t,1)},\cdots,x_{0}^{(t,K)}$ obtained by running a score-based sampler such as ULA on the posterior $p(x_{0}|x_{t})\propto\exp(-E(x_{0})-\lVert x_{t}-\alpha_{t}x_{0}\rVert^{2}/2\sigma_{t}^{2})$. Therefore, in RDMC, obtaining a single sample $x_{t}\sim p(x_{t})$ via ULA at each time step $t\in[1,T]$ involves an intermediate step of generating $K$ posterior samples via another ULA, leading to a nested MCMC sampling procedure that imposes substantial computational demands. In contrast, DiGS does not have such hierarchy since it requires only one MCMC chain during denoising sampling in each Gibbs sweep, significantly reducing the computational burden. Figure 7 compares the computational costs of these two methods when applied to MoG target as shown in Figure 1, demonstrating that DiGS can achieve the same accuracy as RDMC with 10×10$\times$10 × fewer energy evaluations. The efficiency of DiGS is particularly vital in applications such as molecular dynamics (Section 4.3), where even a single energy evaluation is costly.
3.5 Auxiliary Variable MCMC
DiGS belongs to the broader auxiliary variable MCMC family. This family encompasses various notable methods such as the Swendsen-Wang algorithm (Swendsen & Wang, 1987), slice sampling (Neal, 2003), Hamiltonian Monte Carlo (HMC) (Duane et al., 1987; Neal et al., 2011), and auxiliary variational MCMC (Habib & Barber, 2018; Agakov & Barber, 2004), see Barber (2012) for a detailed introduction. Among these, HMC bears the closest resemblance to DiGS. We delve into the similarities and differences in the following discussion.
As introduced in Section 1.1, HMC generates samples from the joint distribution $p(x,v)=p(x)p(v)$ where the $v$ is the auxiliary variable that represents the momentum. The momentum is usually distributed as a Gaussian $p(v)=\mathcal{N}(v|0,\sigma_{v}^{2}I)$ that is independent of $x$. However, there are other variants of HMC where the momentum auxiliary variable $v$ depends on $x$. For example, in Riemannian Manifold HMC (Girolami & Calderhead, 2011), the momentum is distributed as $p(v|x)=\mathcal{N}(v|0,\Sigma_{v}(x))$, where $\Sigma_{v}(x)$ is the Fisher information matrix that captures the local curvature of the energy around $x$. The auxiliary variable in DiGS is the convolved variable $\tilde{x}$, which is a noisy version of $x$, given by $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$. Notably, when $\alpha=0$, it follows that $p(\tilde{x}|x)=p(\tilde{x})=\mathcal{N}(\tilde{x}|0,\sigma^{2}I)$, and the joint distribution $p(x,\tilde{x})=p(x)p(\tilde{x})$ in DiGS is identical to $p(x,v)$ in the classic HMC formulation. Despite this, DiGS employs Gibbs sampling which alternately samples from the two conditionals $p(\tilde{x}|x)$ and $p(x|\tilde{x})$, whereas HMC simulates the Hamiltonian equations as discussed in Section 1.1, interleaved with samples from $p(v)$.

### v5

@@ 2.3 Choosing the Gaussian Convolution Kernels
The performance of DiGS can be influenced by the hyperparameters $\alpha$ and $\sigma$ in the Gaussian convolution kernel. Intuitively, for a given fixed $\alpha$, a large $\sigma$ will enhance the exploration capability of the sampler. However, this also makes the denoising posterior $p(x|\tilde{x})$ closer to $p(x)$ as illustrated in Equation 11, thereby elevating the complexity of denoising sampling. Similarly, with a fixed $\sigma$, reducing $\alpha$ will bring the modes closer but will also make $\nabla_{x}\log p(x|\tilde{x})$ close to $\nabla_{x}\log p(x)$ as shown in Equation 12, making the denoising sampling challenging. To illustrate the effects of hyperparameters, we apply DiGS to the MoG problem described in Figure 1 with varying values of $\alpha,\sigma$ and show the results in Figures 4(a) and 4(b), respectively. This demonstrates that within a specific range of $\alpha$ and $\sigma$ values, DiGS consistently achieves optimal performance indicated by almost zero MMD. However, the quality of samples degrades when hyperparameters deviate beyond certain thresholds.
Although these results suggest that DiGS is robust to a certain range of hyperparameters, selecting an appropriate range remains crucial for each specific target distribution. The optimal range can depend on the support and the shape of the target density, which is typically unknown in practice. Next section will introduce a multi-level noise scheduling to mitigate the need for precise hyperparameter selection.
@@ 2.4 Multi-Level Noise Scheduling
Drawing inspiration from the noise scheduling technique used in diffusion models (Ho et al., 2020; Song & Ermon, 2019; Song et al., 2021; Gao et al., 2018), we propose to use a sequence of Gaussian convolution kernels with $10<\alpha_{T}<\cdots<\alpha_{1}<$ 1 and the corresponding variance $\sigma^{2}_{t}=1-\alpha_{t}^{2}$ in each Gaussian convolution kernel $p_{t}(\tilde{x}_{t}|x)=\mathcal{N}(\tilde{x}|\alpha_{t}x,\sigma_{t}^{2}I)$. This approach is commonly known as the variance-preserving (VP) schedule, where $p_{t}(\tilde{x}_{t})=\int p_{t}(\tilde{x}_{t}|x)p(x)\mathop{}\!\mathrm{d}{x}$ and $p_{0}(\tilde{x}_{0})\equiv p(x)$. As $\alpha_{T}\rightarrow 0$, it follows that $p_{T}(\tilde{x}_{T}|x)\rightarrow\mathcal{N}(\tilde{x}_{T}|0,I)$ and $p_{T}(\tilde{x}_{T})\rightarrow\mathcal{N}(\tilde{x}_{T}|0,I)$, which are easy to sample from since they are independent of the characteristics of the target distribution $p(x)$.
We follow the common sampling procedure in the score-based diffusion model literature (Song & Ermon, 2019, 2020) and propose the following sampling procedure: for each $t$ from $T$ to $1$, we run DiGS to generate a sample, which is used as an initial point in the subsequent time step $t{-}1$. For a given number of noise levels $T$, we apply a simple linear scheduling scheme to determine the values of $\alpha_{t}$. Specifically, given the end points $\alpha_{1}$ and $\alpha_{T}$, we have
$\alpha_{t}=\alpha_{T}+(\alpha_{1}-\alpha_{T})\frac{T-t}{T-1},\quad\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}.$ (17)
@@ 3 Comparison to Related Methods
Figure 5: Comparison of a multi-modal target distribution $p(x)$, tempered distribution $p_{\beta}(x)$, and convolved distribution $p(\tilde{x})$.
@@ 3.1 Tempering-Based Sampling
where $\beta\equiv 1/\tau<1$ is the inverse temperature. As $\tau\rightarrow\infty$, the tempered target $p_{\beta\rightarrow 0}(x)$ converges to a flat distribution, which encourages transitions among different modes. Tempering is a key building block to develop state-of-the-art multi-modal sampling methods such as parallel tempering (PT) (Swendsen & Wang, 1986; Geyer & Thompson, 1995; Syed et al., 2022; Surjanovic et al., 2022) and annealed importance sampling (AIS) (Neal, 2001).
@@
To empirically compare the mode-bridging property of PT and DiGS, we conduct experiments on an extreme problem, “Mixture of Deltas”, which is an MoG with extremely small variance in each Gaussian. Figure 6 shows the comparison of the samples from these two samplers with the same budget: PT is unable to escape the central mode in this extreme setting even with large temperatures $\tau$, whereas DiGS manages to recover all 9 modes, demonstrating its superiority over PT in terms of mode exploration and coverage. To make PT work on this problem, one needs to use more chains with the DEO scheme (Deng et al., 2023); see Appendix D.4 for a detailed description.
@@
Figure 6: Comparison between parallel tempering (PT) and DiGS on a “Mixture of Deltas” problem in 2D with the same budge. PT consists of 5 temperatures, while DiGS uses one noise level. Each sampler is initialized at the origin and generates 1,000 samples. Detailed experimental setup can be found in Appendix D.4.
@@ 3.2 Score-Based Diffusion Model
The convolution technique plays a key role in score-based diffusion models (Song & Ermon, 2019; Song et al., 2021; Song & Ermon, 2020; Ho et al., 2020; Sohl-Dickstein et al., 2015; Gao et al., 2020). These models employ a series of convolutions (indexed by $t$) to form a distribution sequence from $p(\tilde{x}_{0})\equiv p(x)$ to a simple distribution $p(\tilde{x}_{T})$. At each time step $t$, the score function at time $\nabla_{\tilde{x}_{t}}\log p(\tilde{x}_{t})$ is learned from samples ${x_{1},\cdots,x_{N}}\sim p(x)$ using denoising score matching (Vincent, 2011). To sample from $p(x)$, ULA and Euler discretization are typically applied in reverse ($t=T\to 0$), using samples from each time step $t$ as the initial point at time $t{-}1$. This scheme has shown state-of-the-art generation quality for complex data like images. However, unlike score-based/diffusion generative models where $\nabla_{\tilde{x}}\log p(\tilde{x})$ is learned directly from samples, our setting only assumes access to an energy function $E(x)$ without samples, making the noisy marginal score $\nabla_{\tilde{x}}\log p(\tilde{x})$ intractable, as shown in Equation 9.
@@
Zhang et al. (2023a) proposes a pseudo-Gibbs sampler for the joint $p(x,\tilde{x})=p(x)p(\tilde{x}|x)$ to obtain clean samples from score-based diffusion models, where $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$ and $p(x|\tilde{x})$ is approximated by a full-covariance Gaussian. This approach also requires training data to estimate the score function $\nabla_{\tilde{x}}\log p(\tilde{x})$, which is different from our problem setting.
3.3 Proximal Sampler
@@
DiGS also belong to the proximal sampler family (Chen et al., 2022; Lee et al., 2021), where Gibbs sampling is executed between the Gaussian convolution distribution $p(\tilde{x}|x)$ and the denoising posterior $p(x|\tilde{x})$. In the denoising sampling step, proximal samplers find a mode using gradient-based optimization technique and perform rejection sampling around that mode. We highlight several significant improvements of DiGS over the proximal samplers. First, we employ a score-based sampler such as MALA and HMC in the denoising step instead of rejection sampling around a single mode as in proximal samplers, which helps the method scale to higher-dimensional distributions. Second, we employ an innovative Metropolis-within-Gibbs scheme to initialize the denoising sampling step, demonstrating its importance in achieving correct density allocation in multi-modal situations, as discussed in Section 2.2. Third, we emphasize the importance of tuning Gaussian convolution kernel in Section 2.4 and employ a multi-level scheduling to eliminate the need of choosing Gaussian convolution hyperparameters. These improvements make the DiGS applicable to practical problems with multi-modal target distributions in high dimensional spaces.
@@
Entropy-MCMC (Li & Zhang, 2023) is designed to sample from the flat regions within the posterior of the Bayesian neural network $p(\theta|\mathcal{D})$, where $\mathcal{D}$ represents the training dataset. Specifically, the approach begins by defining a surrogate distribution using Gaussian convolution:
@@
where $p(\theta_{a}|\theta)\propto\exp\left(-\frac{1}{2\eta}\|\theta_{a}-\theta\|^{2}_{2}\right)$ is a Gaussian kernel aimed at smoothing out the sharp regions in the posterior. Subsequently, SGLD (Welling & Teh, 2011) is applied to sample within the joint space of $(\theta,\theta_{a})$, where samples of both $\theta$ and $\theta_{a}$ are kept. It is crucial to highlight that, while this method aims to seek flat modes in the target distribution $p(\theta|\mathcal{D})$, similar to the approaches described in Chaudhari et al. (2019); Staines & Barber (2012), it is not intended and cannot provide exact samples from the target distribution $p(\theta|\mathcal{D})$ due to its optimization nature. In contrast, DiGS is designed as a valid MCMC sampler that is specifically designed to identify all modes (including both sharp and flat modes) and their corresponding density allocations in the target distribution, thus pursuing a distinct objective.
Figure 7: Computational cost comparison between DiGS and RDMC. The x-axis is the number of energy evaluations (×$\times 10^{6}$) and the y-axis is the MMD between true and generated samples. We use DiGS with a single noise level $(\alpha=1,\sigma=1)$ and run Gibbs sampling for 1-10 sweeps, represented by the 10 blue scatters. The RDMC is experimented with $T\in\{1,2,3,4\}$, represented by the four orange points in the plot. Detailed experimental setup can be found in Appendix D.5.
@@
3.4 Reverse Diffusion Monte Carlo
$\nabla_{x_{t}}\log p(x_{t})$ $=\left(\alpha\mu(x_{t})-x_{t}\right)/\sigma_{t}^{2},$ (20)
where the denoising posterior mean $\mu(x_{t})=\int xp(x|x_{t})\approx\frac{1}{K}\sum_{k=1}^{K}x_{0}^{(t,k)}$ is approximated by the Monte Carlo method, with samples $x_{0}^{(t,1)},\cdots,x_{0}^{(t,K)}$ obtained by running a score-based sampler such as ULA targeting the posterior $p(x_{0}|x_{t})\propto\exp(-E(x_{0})-\lVert x_{t}-\alpha_{t}x_{0}\rVert^{2}/2\sigma_{t}^{2})$. Therefore, in RDMC, obtaining a single sample $x_{t}\sim p(x_{t})$ via ULA at each time step $t\in[1,T]$ involves an intermediate step of generating $K$ posterior samples via another ULA, leading to a nested MCMC sampling procedure that imposes substantial computational demands. In contrast, DiGS does not have such hierarchy since it requires only one MCMC chain during the denoising sampling step in each Gibbs sweep, significantly reducing the computational burden. Figure 7 compares the computational costs of these two methods when applied to MoG target as shown in Figure 1, demonstrating that DiGS can achieve the same accuracy as RDMC with 10×10$\times$10 × fewer energy evaluations. The efficiency of DiGS is particularly vital in applications such as molecular configuration sampling (Section 4.3), where even a single energy evaluation is costly.
3.5 Auxiliary Variable MCMC
DiGS belongs to the broader auxiliary variable MCMC family. This family encompasses various notable methods such as the Swendsen-Wang algorithm (Swendsen & Wang, 1987), slice sampling (Neal, 2003), Hamiltonian Monte Carlo (HMC) (Duane et al., 1987; Neal et al., 2011), and auxiliary variational MCMC (Habib & Barber, 2018; Agakov & Barber, 2004); see Barber (2012) for a detailed introduction. Among these, HMC bears the closest resemblance to DiGS. We delve into the similarities and differences below.
As introduced in Section 1.1, HMC generates samples from the joint distribution $p(x,v)=p(x)p(v)$ where the $v$ is the auxiliary variable that represents the momentum. The momentum is usually distributed as a Gaussian $p(v)=\mathcal{N}(v|0,\sigma_{v}^{2}I)$ that is independent of $x$. However, there are other variants of HMC where the momentum auxiliary variable $v$ depends on $x$. For example, in Riemannian Manifold HMC (Girolami & Calderhead, 2011), the momentum is distributed as $p(v|x)=\mathcal{N}(v|0,\Sigma_{v}(x))$, where $\Sigma_{v}(x)$ is the Fisher information matrix that captures the local curvature of the energy around $x$. The auxiliary variable in DiGS is the convolved variable $\tilde{x}$, which is a noisy version of $x$, given by $p(\tilde{x}|x)=\mathcal{N}(\tilde{x}|\alpha x,\sigma^{2}I)$. Notably, when $\alpha=0$, it follows that $p(\tilde{x}|x)=p(\tilde{x})=\mathcal{N}(\tilde{x}|0,\sigma^{2}I)$, and the joint distribution $p(x,\tilde{x})=p(x)p(\tilde{x})$ in DiGS is identical to $p(x,v)$ in the classic HMC formulation. Despite this similarity, DiGS employs Gibbs sampling which alternately samples from the two conditionals $p(\tilde{x}|x)$ and $p(x|\tilde{x})$, whereas HMC simulates the Hamiltonian equations as discussed in Section 1.1, interleaved with samples from $p(v)$.